OpenStudy (anonymous):

Which statement is best represented by ? 1/4*>(with the line under it)12 One fourth times a number is less than 12. One fourth times a number is at most 12. One fourth times a number is greater than 12. One fourth times a number is at least 12.

Nnesha (nnesha):

\[\frac{ 1 }{ 4} \ge 12\]

OpenStudy (anonymous):

thats what it looks like

Nnesha (nnesha):

\[\ge \] greater than , at least \[\le \] less than or equal to , at most

Nnesha (nnesha):

there are definition now what you think what is your answer

OpenStudy (anonymous):

1/4* a number that is greater than or equal to 12 as i learned it

OpenStudy (anonymous):

I thought it was One fourth times a number is greater than 12. but thats wrong

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thank you

Nnesha (nnesha):

yes that like i said \[\ge \] this is greater than so ys

OpenStudy (anonymous):

dont worry you tried and thats what counts

OpenStudy (anonymous):

lol thank you i got the signs mixed up :)

OpenStudy (anonymous):

always ready to help

Nnesha (nnesha):

so what is your answer ??

OpenStudy (anonymous):

One fourth times a number is at least 12.

Nnesha (nnesha):

@daphne101 yeah that's good welcome to open study http://openstudy.com/code-of-conduct <--- hope you read this where direct answer not allowed

OpenStudy (anonymous):

are you actually a red head or is that just the un just curious sorry if im getting too personal and rude

Nnesha (nnesha):

who are talking to ???

Nnesha (nnesha):

you**

OpenStudy (anonymous):

sorry that was my sis

OpenStudy (anonymous):

shes very rude and obnoxiouss

Nnesha (nnesha):

srsly ?? :O its okay :)